OpenStudy (anonymous):

a recreational swimming lake is treated to control bacteria. suppose t days after a treatment the concentration of bacteria per cubic centimeter is given by c(t)=30t^2-240t+500; 0

OpenStudy (anonymous):

differentiate 60t - 240 = 0 for minmm t = 6 days minm concentrtation is 30(36) - 240*6 + 500 = 1080 - 1440 + 500 = 140

OpenStudy (anonymous):

oh sorryyyyy - i cant count! t = 4 days! minm concentration = 30*16 - 240*4 + 500

OpenStudy (anonymous):

do u follow that ok ?

OpenStudy (anonymous):

my brain did a 'switch off' for a moment

OpenStudy (anonymous):

yeah i think! so the minimum concentration would be 20?

OpenStudy (anonymous):

480 -960 + 500 = 20 right

OpenStudy (anonymous):

awesome thanks so much!!!

OpenStudy (anonymous):

no probs